find the point of intersection of the tangent drawn  to the curve x2y=1-y at the point where it is intersected by the curve xy=1-y

Grade:12

1 Answers

Dharma Teja
7 years ago
given x2y=1-y and xy=1-y
solving by taking x2y =xy
we get two points (1,1/2) and (0,1)
now differentiate x2y=1-y
and get slopes for both the points
get tangent equation by formula Y-y=m(X-x)
we get tangents equations as y=1/2 and x-y+1=0
solving these point of intersection is (-1/2,1/2)